OpenStudy (sleepyjess):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 4 * 6 + 5 * 7 + 6 * 8 + ... + 4n( 4n + 2) = 4(4n+1)(8n+7)/6

OpenStudy (sleepyjess):

Here is the end fraction \(\dfrac{4(4n+1)(8n+7)}6\)

sammixboo (sammixboo):

But.. I can't.. Jess D: @Luigi0210

OpenStudy (sleepyjess):

Sammi... tsk tsk tsk I was counting on you :(

OpenStudy (sleepyjess):

From what I can tell from the lessons, I just replace n with 1 and see if it equals 4, n with 2 see if it is 6, etc

OpenStudy (luigi0210):

I'm trying to remember how to do this.. this stuff is old to me xD

OpenStudy (luigi0210):

Yea, the first thing to do would be to plug in n=1 to see if the statement is true.

OpenStudy (sleepyjess):

ok, I got 60 for n = 1, but if it is true, is it supposed to equal 4 or 24?

OpenStudy (anonymous):

\[\frac{4(4 \times 1+1)(8(1)+7)}{6}=\frac{4 \times 5 \times 15}{6}=50\] Can't remember how to do left side for n=1

OpenStudy (sleepyjess):

ah whoops, I must have gotten confused haha, yes it would be 50 not 60 :)

OpenStudy (sleepyjess):

@ganeshie8 could you please help?

ganeshie8 (ganeshie8):

for \(n=1\) left hand side contains just the first term : \(4\cdot 6\)

OpenStudy (sleepyjess):

So if it were true, then for n = 1 on the right, it should equal to 24?

ganeshie8 (ganeshie8):

Yes clearly the given statement is false for \(n=1\). So we're done.

OpenStudy (sleepyjess):

Ok, thank you :)

ganeshie8 (ganeshie8):

yw:)